OpenStudy (anonymous):

How many grams of potassium chloride are in 6.90 moles?

OpenStudy (abb0t):

You should be able to recognize the molecular formula for potassium chloride. Kknowing it's molecular formula, you can calculate it's grams by adding up its molar mass. The relationship which is \[\frac{ grams }{ mole }\] so you take: mass of potassium + mass of chlorine = ________ g/ mol Next, use dimensional analysis to cross cancel (similar to how you did in algebra and fractions). For instance: \[\frac{ a }{ 2b } \times \frac{ 4b }{ a } = 2b\] Do you see how that is. It follows a similar method. So do that to cross cancel units in this case.

OpenStudy (anonymous):

by following the above method answer is 514.05 grams , try it your self also

OpenStudy (abb0t):

That's correct @rosewhite but I think you should of let him/her try the problem out rather than just give them the answer. It's a learning process.

OpenStudy (anonymous):

atomic mass of potassium=39 u, atomic mass of chlorine=35.5 u, so the mass of KCl=35.5+39=74.5, let there be x grams of KCl in 6.90 moles so we have \[\huge\frac{ x }{ 74.5 }=6.90=>x=74.5*6.90=514.05 \]

OpenStudy (anonymous):

ok, I will keep that @abb0t